OpenStudy (anonymous):

Find the first five non-zero terms of power series representation centered at x=0 for the function below: 7/(1-x^2) I get that the first term in the series is 7, but I have no clue how to find the other terms. Do I derive the equation? Help.

OpenStudy (anonymous):

$$\frac1{1-x}=1+x+x^2+x^3+x^4+\dots\\\frac1{1-x^2}=1+x^2+x^4+x^6+x^8+\dots\\\frac7{1-x^2}=7+7x^2+7x^4+7x^6+7x^8\dots$$

OpenStudy (anonymous):

Can I ask how you did it? do you just convert to power series and solve?

OpenStudy (anonymous):

well, I used knowledge of geometric series to write it in this form. you can also do it using Taylor's theorem

OpenStudy (amistre64):

you can also just do the division ... -------- 1 - x^2 | 7